### v2

@@ 4 Successful energy transitions rely on balancing trade-offs between local priorities
Figure 5: Obtained values from participants depending on whether a metric was prioritised; larger difference between participants who prioritised a certain metric might hint at more available options to achieve this or at higher weights of this metric. A lower value on the y-axis indicate a lower impact, which is generally desired.
@@
Figure 6: Obtained values for vulnerability and emissions from participants’ choices. Each point in a) represents a response and the according system’s vulnerability and emissions. The dashed outer line represents the complete near-optimal feasible space and the colour of a point indicates the resulting slack. The convex hulls for seven slack levels in b) show how the decision space changes with increasing slack.
As participants tried to balance multiple priorities, the evaluation of how well people manage to achieve their priorities is complicated (given that their individual weighing of the different metrics is unknown to us). Figure 6a), illustrates how participants balanced the two most commonly prioritised metrics, vulnerability and emissions, and the resulting slack value through the shade of the point. Most of the participants have chosen systems with very low emissions with larger variability for vulnerability. Furthermore, figure 6b) shows how the decision space of the participants grows with increasing slack values. While it is possible to balance these two priorities, even at low slack levels, further reductions of vulnerability come at the expense of a more expensive system.
Figure 7: Correlation matrix between revealed slider choice and stated priority of participants.
@@
Table 1: Exemplary qualitative feedback from the participants. They are later referred to in text through their feedback ID.
Feedback ID Quote
R02 ”Interesting interlinked scenario. Funny to play around with different parameters and see how complex the energy mix is and consequences are.”
R03 “Well done and gets you thinking.”
R04 ”Cannot choose desired energy solution (i.e. 100% return to coal)”
R05 ”With respect to the current situation, a coal-fired power plant, which was already in operation, would be the most sensible economic and environmentally friendly solution. Longyearbyen is not equipped for the green transition”
R07 “Look at nuclear power. Safe, space-effective and durable.”
[…]
00footnotetext: The full list of qualitative feedback can be found in the supplementary material.
@@
5 Discussion and conclusion
We use the near-optimal feasible space of an energy system model to generate 56,050 alternative energy system designs for Longyearbyen, Svalbard. These system configurations were presented to 126 local citizens who explored different investments in energy system components through an interactive interface. It allowed the participants to explore trade-offs between and implications of different options. Compared to previous interactive modelling tools [18, 21], the deeper integration through near-optimal modelling incorporated system interactions of different components as well as intermittency of renewables. Most critically, this approach limits the decision space to feasible system configurations. This prevents users from selecting systems that fail to reliably supply energy, which is of utmost importance in vulnerable communities like Longyearbyen.
@@ 5 Discussion and conclusion
The five dimensions/sliders that the participants were adjusting were meant to represent the most relevant technologies, following the local energy transition plan [41]. Although the modelling reflected political and technological realities, both during the process and in the post-study feedback, participants mentioned technologies such as coal, nuclear and geothermal energy that did not meet these criteria. Mentions of returning to the recently discontinued coal-fired thermal power plant can be tied to Longyearbyen’s history as a coal mining settlement which formed local identity and legitimised Norwegian presence on Svalbard [42]. The interest in technologies that are considered infeasible (see Methods as to why) highlights the importance of participatory studies as they reveal opinions, and help to understand why they are formed and keep society informed. Furthermore, some participants stated a preference for reduced demand / consumption, which could indeed reduce trade-offs and repercussions of new energy infrastructure developments. This would be an interesting dimension to include in such an interface, but how it is presented needs to be carefully considered: from a system perspective it may appear desirable, but societal and individual impacts would need to be illustrated effectively.
The interface and model are highly customisable and many insights are transferable to other circumstances; some challenges remain and certain results may be specific to Longyearbyen. Participants in Longyearbyen are expected to be relatively well-informed: not only is the energy system small and familiar, its transition has also been publicly discussed for many years. The remoteness and bitterly cold climate together with geopolitical tensions expose risks to various types of interruptions, and a greater degree of vulnerability than elsewhere. Besides, most local communities might not directly experience all impacts of their decisions, e.g. if they are connected to neighbouring areas and can import energy. Moreover, future applications should consider engaging with participants who spend more time with the interface to enhance the understanding of the processes and methods. Whereas we achieved a good number of participants and good diversity in terms of demographics, through the interactions in public spaces, this limited the time available for participants to get acquainted with the interface and its workings. A few users were concerned about the complexity of the underlying modelling, whereas others stated that it was insightful and enjoyable and suggested making the interface available on the community council’s website. It should be noted, however, that no participant, to our knowledge, utilised or read the information sheet that we prepared and laid out. Instead of a full overview of the modelling assumptions, participants preferred direct enquiries for specific explanations. Future work should also look to include stakeholders throughout the processes, and not only at a single stage, as with our particular application. A greater degree of participation may be achieved by, for example, co-designing the dimensions (decision-variables) and relevant metrics together with stakeholders.
Furthermore, the slack or admissible cost increase in near-optimal workflows impacts which energy system configurations are possible. In contrast to common values of up to 10-15% [43, 44], we purposefully increase slack levels (125%) to broaden the diversity of options, and instead let participants decide the slack level of their system. The stated willingness-to-pay from the post-study questionnaire has a median of 15%, similar to both previous studies and within the range of how close to cost-optimal previous transitions have been (9-23%) [45]. On the other hand, the revealed willingness-to-pay from the submitted energy system designs has a median of 91%. This indicates tensions between preconceptions on “acceptable” deviations from cost-optimality and the realisation of their own preferences. Further research into this topic may reveal realistic slack values for future near-optimal modelling.
@@
Growing interest [25, 34] and recent advances of near-optimal methods in energy system modelling provide opportunities for further fine-tuning of the presented methodology. For example, an MGA post-processing by Lau et al. [46] can respond “live” to stakeholder input when exploring the near-optimal space, potentially reducing pre-computations as well as modelling assumptions. Lombardi and Pfenninger [35] propose a Human-in-the-loop (HITL) extension to conventional MGA approaches, which generates a secondary collection of near-optimal alternatives closely aligned with initially selected preferred solutions. These, and other improvements to MGA and near-optimality in energy system models may contribute to further enhancing its usefulness for stakeholder interactions, elicitation of informed stakeholder preferences and as a tool for learning.
@@
For the energy transition to succeed, social and political feasibility is of utmost importance. Contested technologies and infrastructure risk slowing down processes and limit political will. Supporting legitimacy in decisions and processes, and making people’s voices heard in energy transitions improves our chance of collectively limiting climate change and staying within the planetary boundaries. We have, with this work, demonstrated a methodology for combining near-optimality with stakeholder interactions that shows promise for illustrating and navigating difficult trade-offs in energy systems planning. Participants who have interacted with the methodology have clearly shown that there is more to consider than only cost-optimality, and that all preferences cannot always be met. Engaging with a continuum of feasible system designs via visualisation tools, such as our interactive interface, can help demonstrate the complexity in planning decisions, and also create meaning and a positive feeling around the challenges and dilemmas of energy transitions.
@@ 6 Methods
This study is based on an interface for exploring near-optimal solutions to an optimisation model for the Longyearbyen energy system. By introducing the possibility of exploring near-optimal solutions (defined below in more detail) we can assess how different stakeholders prioritise various aspects of the energy transition and the impact on their community. For example, one might be willing to have a slightly more expensive system to minimise a technology (e.g. solar power) or maximise other socially relevant metrics [25]. Our results are mainly based on preferences submitted by study participants through said interface, which presents the participants with sliders controlling central design parameters and a display showing the impact of design choices on a number of different metrics.
@@
The model used in this study, PyPSA-Longyearbyen or PyPSA-LYB, is an energy system optimisation model (based on the open-source modelling framework PyPSA), which optimises investment and operational decisions for a future energy system in Longyearbyen. This model was commissioned by Longyearbyen Lokalstyret (the community council in Longyearbyen) and Svalbard Energi AS and developed by Multiconsult for the report “Energiplan Longyearbyen” [41] in 2023. A number of different scenarios were explored in said report; for this study we select the model corresponding to a baseline scenario with a 2000 NOK/tCO2 carbon price. We provide a description of the model for completeness (below), but do not make any modifications to the model as obtained from Svalbard Energi.
@@
Coal power was excluded from the model and interface based on the political decisions that the energy supply in Longyearbyen should support and be in line with Norway’s climate targets for 2030 and 2050 [47]. Nuclear is not considered techno-economically feasible, as both conventional nuclear and small modular reactors would be over-dimensioned in Longyearbyen.
@@
A near-optimal solution (strictly speaking an $\varepsilon$-near-optimal feasible solution) describes an alternative, feasible system design whose costs are $\varepsilon$ higher than in the cost-optimal solution. Such near-optimal solutions can be desirable because they might outperform the cost-minimal solution with regards to other qualities. Recent research on near-optimal solutions for energy systems has shown that these system configurations can vary widely without large cost differences [32, 31]. Showcasing many near-optimal solutions can help stakeholders understand the flexibility in planning of systems; if we describe near-optimal spaces geometrically, we can furthermore systematically map out trade-offs.
@@
The space of near-optimal solutions is a convex polytope, being the feasible space of a linear program (LP), but typically consists of too many vertices in too many dimensions to be directly useful. Thus, we follow Grochowicz et al. [32] by approximating a projection onto certain key dimensions of the near-optimal feasible space; the approximation with a limited number of vertices is obtained by repeatedly solving the original model LP with a near-optimality constraint using different, random objective functions.
In this study, the key dimensions we project down to are precisely the variables that participants are allowed to control with sliders: total investment in wind power, solar, heat storage (both overground water tank and geothermal), hydrogen storage (including storage tanks, electrolysers and fuel cells) as well as the total cost of green fuel imports (including ammonia, biogas and pellets).
@@
We chose a maximum slack level of 125% to allow vastly different system configurations; country-level studies often use lower values instead, with one hindcasting study concluding in historic deviations from the cost-optimum of 9–24% in the UK [45]). The method presented in Grochowicz et al. [32] leads to an approximation of the near-optimal space by a specified number of boundary vertices — the above paper using 150 optimisations. For the interpolation of metrics to work well, this study is based instead on a total of 56,050 optimisations at a number of different slack levels between 0 and 125%. This way, we obtain near-optimal solutions (for which metrics can be computed) throughout the interior and not only on the boundary of the near-optimal space.
@@
    where $c$ is 1 - the share of VRE, $d$ is 1 - share of imported energy, $e$ is 1 - the share of different technologies present in the system at an installed capacity of at least 1 MW, and $f=\min{1,heat_{pnorm}}$ with $heat_{pnorm}$ the normalised capacity of heat storage (a value of $1$, corresponding to $4$ GWh, can cover heat demand for two weeks in the winter, and one month in the summer). In effect, a higher share of VRE, imported energy, as well as a higher complexity of the system increase vulnerability (the value of $V$), whereas higher heat storage reduces vulnerability.
@@
Table 3: Assumed land use requirements for technologies included in the modelling (either explicitly in the sliders or implicitly in the model). The land use requirement of different technologies is based on Chen et al. [48] and Sasse et al. [49].
@@
The interface presented to study participants consists of two elements (figure 3): sliders allowing the user to select and tweak an energy system design, and a display of a number of metrics associated with the selected system design (see above).
Importantly, the sliders only allow participants to select feasible system designs at most 125% more expansive than cost-optimal. Formally, the slider values may only encode the coordinates of points within the near-optimal space of the underlaying energy system model. This is achieved by dynamically limiting the selectable ranges of each slider based on the currently chosen design. Specifically, let $x_{1},x_{2},x_{3},x_{4},x_{5}$ represent currently selected slider values, corresponding to the key dimensions described above (wind, solar, green fuel imports, heat infrastructure and hydrogen storage), and let $\mathcal{F}_{\varepsilon}$ by the projection of the near-optimal space down to these five dimensions, with $\varepsilon=125\%$. Then $(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathcal{F}_{\varepsilon}$. The interface user can change a single variable at a time to select a new system design; without loss of generality, suppose this is $x_{1}$. Allowable new values for $x_{1}$ are exactly those $x_{1}^{*}$ such that $(x_{1}^{*},x_{2},x_{3},x_{4},x_{5})\in\mathcal{F}_{\varepsilon}$. Since $\mathcal{F}_{\varepsilon}$ is convex, this condition is satisfied on a line segment $x_{1}^{*}\in[a,b]$, with
@@
When the user happens to exactly select a system design for which pre-computed metrics were recorded, the interface simply displays those pre-computed values. Since slider values can be changed continuously (up to the limits of pixel and floating point precision) however, it is much more likely that the user selects a system design for which no pre-computed metrics are available. In this case, a heuristic is used to compute approximate values for each of the metrics. The heuristic consists of a linear interpolation among nearby pre-computed metric values. Linear interpolation of metrics has previously been suggested by Lau et al. [46], who treat points in a near-optimal space as linear combinations of all computed near-optimal model solutions. While their method works well with a relatively small sample of near-optimal solutions, we apply a near-neighbour filter first in order to further reduce the computational complexity of the heuristic.
@@
Table 4: Demographic characteristics of the sample, the population of Longyearbyen and Norway as a whole
@@
00footnotetext: Data for Norway is from 2023 and information of their income was optional, leading to a smaller sample.
@@
Author contributions All authors conceived the idea and designed the study. O.V., A.G. and K.vG. did the field work and collected the data. A.G. lead the implementation of the near-optimal modelling. K.vG. lead the development of the interactive interface. O.V. lead the writing of the paper, with considerable input from the other authors.
@@
The field data that support part of the findings of this study are not available owing to the confidentiality of participants’ identities. Longyearbyen is a small settlement, and providing individual data points may risk the identification of particular individuals, even if the data is anonymised. Aggregated data can be made available upon request.

### v3

@@ 4 Successful energy transitions rely on balancing trade-offs between local priorities
Figure 5: Obtained values from participants depending on whether a metric was prioritised, along with associated vulnerability and emissions data; larger difference in a) between participants who prioritised a certain metric might hint at more available options to achieve this or at higher weights of this metric. A lower value on the y-axis indicate a lower impact, which is generally desired (rescaled relative to minimum and maximum values achieved by participants). In the box plot, the central lines represent median values, the box limits represent the upper and lower quartiles, the whiskers extend from the box to the farthest data point lying within 1.5x the IQR. Outliers (fliers) are also shown. The median values as well as first and third quartiles are given in the supplementary information. Each point in b) represents a response and the according system’s vulnerability and emissions. The dashed outer line represents the complete near-optimal feasible space and the colour of a point indicates the resulting slack. The histograms show the frequency of results for emissions (horizontally) and vulnerability (vertically), with 14 bins. The convex hulls for seven slack levels in c) show how the decision space changes with increasing slack.
@@
As participants tried to balance multiple priorities, the evaluation of how well people manage to achieve their priorities is complicated (given that their individual weighing of the different metrics is unknown to us). Figure 5b), illustrates how participants balanced the two most commonly prioritised metrics, vulnerability and emissions, and the resulting slack value through the shade of the point. Most of the participants have chosen systems with very low emissions with larger variability for vulnerability. Furthermore, figure 5c) shows how the decision space of the participants grows with increasing slack values. While it is possible to balance these two priorities, even at low slack levels, further reductions of vulnerability come at the expense of a more expensive system.
Figure 6: Correlation matrix between revealed slider choice and stated priority of participants. Correlations are pairwise and based on the Pearson correlation coefficient (PCC).
@@
5 Discussion
We use the near-optimal feasible space of an energy system model to generate 56,050 different energy system designs for Longyearbyen, Svalbard. These system configurations were presented to 126 local citizens who explored different investments in energy system components through an interactive interface. It allowed the participants to explore trade-offs between and implications of different options. Compared to previous interactive modelling tools [volkenPerspectivesInformedCitizen2018, holzerSwissElectricitySupply2023, duboisInformedCitizenPanels2019, xexakisAreInteractiveWebtools2019, xexakisModelsWrongTrack2020, flackeInteractivePlanningSupport2017, allenCarbonReductionScenarios2013, demskiEffectsExemplarScenarios2017], applying near-optimal modelling for participatory exercises involving stakeholders allows retaining system interactions of different components as well as intermittency of renewables. Most critically, this approach limits the decision space to feasible system configurations. This prevents users from selecting systems that fail to reliably supply energy, which is of utmost importance in vulnerable communities like Longyearbyen. Additionally, this framework opens up for discussing a diversity of options which cost-optimal scenarios overlook, reducing researcher bias, especially if stakeholders are involved at an early stage of the modelling process.
@@ 5 Discussion
The five dimensions/sliders that the participants were adjusting were meant to represent the most relevant technologies, based on the local energy transition plan [grotteEnergiplanLongyearbyenEnergiomstilling2023]: our modelling followed the political and technological realities [grotteEnergiplanLongyearbyenEnergiomstilling2023], yet both during the process and in the post-study feedback, participants mentioned the absence of desired technologies like coal, nuclear and geothermal energy. Despite unprecedented warming and landslides in town, mentions of reviving the discontinued coal-fired thermal power plant relate to Longyearbyen’s coal mining history which formed local identity and legitimised Norwegian presence on Svalbard [nygaardPhaseoutsEdgeWorld2024]. The interest in technologies that are considered infeasible (see the Methods as to why) highlights the importance of participatory studies as they reveal opinions, and help to understand why they are formed and keep society informed. Furthermore, some participants stated a preference for reduced demand / consumption, which could indeed reduce trade-offs and repercussions of new energy infrastructure developments. This would be an interesting dimension to include in such an interface, but how it is presented needs to be carefully considered: from a system perspective it may appear desirable, but societal and individual impacts would need to be illustrated effectively.
Limitations and Future Directions
Participatory modelling aims to involve stakeholders, removing modeller biases in favour of exploring a breadth of socially viable alternatives. While we demonstrate a new way for stakeholders to explore feasible system designs, large in-depth studies could involve stakeholders at additional stages in the participatory modelling workflow (figure 2). This could involve workshops informing participants about realistic options related to policy targets (e.g. decomissioning coal for climate mitigation), but also incorporating stakeholders in co-designing the study and interface (selecting dimensions, technologies, and metrics) directly through surveying, focus groups or other methods [mcgookinAdvancingParticipatoryEnergy2024]. In this demonstration study, we limit our focus to the effectiveness of interface interactions.
The trade-offs around energy transitions are inherently complex, making it challenging for participants to learn from interactive interfaces in short amounts of time. Indeed, interactive interfaces have been shown to not necessarily create more insight and understanding of the underlaying energy system dynamics than the static presentation of a few selected scenarios [xexakisAreInteractiveWebtools2019] in the context of electronic surveys. In direct feedback (table 5), a few users were concerned about the complexity of the underlying modelling, whereas others stated that it was insightful and enjoyable and suggested making the interface available on the community council’s website. We do not compare learning outcomes to alternatives such as a survey including fewer options, which could lead to less information overload but would be of limited utility in exploring diverse preferences. Interactive tools may also suffer from anchoring effects [demskiEffectsExemplarScenarios2017], where the initially selected system design can influence the solution submitted by the participants. While our study used the same (arbitrary) starting point for all participants, future work could consider using different randomly selected starting points for each participant.
Moreover, our study is based on relatively short interactions (typically in the range of 5–20 minutes) in public spaces, which enabled us to capture a broadly representative sample of the population (see Methods), but means that participants did not have time to learn about our modelling setup in depth. Participants in Longyearbyen are expected to be relatively well-informed: not only is the energy system small and familiar, its transition has also been publicly discussed for many years. The bitterly cold climate together with geopolitical tensions and impacts of climate change expose risks to various types of interruptions, and a greater degree of vulnerability than elsewhere. It should be noted that no participant, to our knowledge, utilised or read the information sheet that we prepared and laid out. Instead, participants preferred and frequently made use of direct enquiries for specific explanations — one of the supervising researchers was always at hand to answer questions.
The discrepancy between stated willingness-to-pay (median of 15% increase over optimality) and actual cost of submitted system designs (median of 91% increase over optimality) highlights some of the complexities of participatory modelling. While the stated willingness-to-pay is similar to both previous studies and common MGA assumptions [trutnevyteDoesCostOptimization2016, lombardiNearoptimalEnergyPlanning2025], even the 91% cost increase over optimality obtained from the participants lies within the range of observed deviations from the cost-optimum [fossastenasEmpiricalEvidenceSlack2025]. In effect, participants seem willing to choose expensive solutions in order to pursue non-monetary preferences, but when asked directly, are not willing to front the cost themselves. Future studies could look into ways of making total system cost more tangible to the interface users (for instance, framing it directly as a personal cost or expected utilities bill) in order to close the gap. In the specific case of Longyearbyen, however, some study participants may have plausibly believed that they would not personally pay for an expensive system, since the whole community is already heavily subsidised by the Norwegian government. Thus, the large gap between stated and revealed system cost slack may be a location-specific artifact to an extent. Further research into this topic may reveal realistic slack values for future near-optimal modelling.
In spite of the limitations outlined above, we show that participants were able to meaningfully pursue their stated objectives (figure 5), indicating that the proposed interactive interface is effective. Still, more in-depth engagements such as focus groups or workshops [mcgookinAdvancingParticipatoryEnergy2024] could lead to even more informed responses (though at higher cost).
@@
Growing interest [vageroCanWeOptimise2023, lauMeasuringExplorationEvaluation2024] and recent advances of near-optimal methods in energy system modelling provide opportunities for further fine-tuning of the presented methodology. For example, an MGA post-processing by Lau et al. [lauModellingGenerateContinuous2024] can respond “live” to stakeholder input when exploring the near-optimal space, potentially reducing pre-computations as well as modelling assumptions. Lombardi and Pfenninger [lombardiHumanintheloopMGAGenerate2025] propose a Human-in-the-loop (HITL) extension to conventional MGA approaches, which generates a secondary collection of near-optimal alternatives closely aligned with initially selected preferred solutions. These, and other improvements to MGA and near-optimality in energy system models may contribute to further enhancing its usefulness for stakeholder interactions, elicitation of informed stakeholder preferences and as a tool for learning.
Conclusion
@@
For the energy transition to succeed, social and political feasibility is of utmost importance. Contested technologies and infrastructure risk slowing down processes and limit political will. Supporting legitimacy in decisions and processes, and making people’s voices heard in energy transitions improves our chance of collectively limiting climate change and staying within the planetary boundaries. We have, with this work, demonstrated a methodology for combining near-optimality with stakeholder interactions that shows promise for illustrating and navigating difficult trade-offs in energy systems planning. Participants who have interacted with the methodology have clearly shown that there is more to consider than only cost-optimality, and that all preferences cannot always be met. Engaging with a continuum of feasible system designs via visualisation tools, such as our interactive interface, can help demonstrate the complexity in planning decisions (one of the key issues of participatory modelling) [mcgookinAdvancingParticipatoryEnergy2024], and also create meaning and a positive feeling around the challenges and dilemmas of energy transitions.
@@ 6 Methods
This study is based on an interface for exploring near-optimal solutions to an optimisation model for the Longyearbyen energy system. By introducing the possibility of exploring near-optimal solutions (defined below in more detail) we can assess how different stakeholders prioritise various aspects of the energy transition and the impact on their community. For example, one might be willing to have a slightly more expensive system to minimise a technology (e.g. solar power) or maximise other socially relevant metrics [vageroCanWeOptimise2023]. Our results are mainly based on preferences submitted by study participants through said interface, which presents the participants with sliders controlling central design parameters and a display showing the impact of design choices on a number of different metrics.
@@
The model used in this study, PyPSA-Longyearbyen or PyPSA-LYB, is an energy system optimisation model (based on the open-source modelling framework PyPSA), which optimises investment and operational decisions for a future energy system in Longyearbyen. This model was commissioned by Longyearbyen Lokalstyret (the community council in Longyearbyen) and Svalbard Energi AS and developed by Multiconsult for the report “Energiplan Longyearbyen” [grotteEnergiplanLongyearbyenEnergiomstilling2023] in 2023. A number of different scenarios were explored in said report; for this study we select the model corresponding to a baseline scenario with a 2000 NOK/tCO2 carbon price. We provide a description of the model for completeness (below), but do not make any modifications to the model as obtained from Svalbard Energi.
The model covers the full energy system of Longyearbyen, including electricity, heat, as well as fully electrified transport and industrial sectors. Although the current transportation system is almost fully based on fossil fuels, many of the large organisations in Longyearbyen have implemented policies for transitioning to an electrified vehicle fleet. With additional policy measures, it is considered possible to have an electrified transportation sector by 2030 [grotteEnergiplanLongyearbyenEnergiomstilling2023].
@@
Coal power was excluded from the model and interface based on the political decisions that the energy supply in Longyearbyen should support and be in line with Norway’s climate targets for 2030 and 2050 [thenorwegianministryofjusticeandpublicsecurityMeldSt262024]. Nuclear is not considered techno-economically feasible, as both conventional nuclear and small modular reactors would be over-dimensioned in Longyearbyen.
@@
A near-optimal solution (strictly speaking an $\varepsilon$-near-optimal feasible solution) describes an alternative, feasible system design whose costs are $\varepsilon$ higher than in the cost-optimal solution. Such near-optimal solutions can be desirable because they might outperform the cost-minimal solution with regards to other qualities. Recent research on near-optimal solutions for energy systems has shown that these system configurations can vary widely without large cost differences [grochowiczIntersectingNearoptimalSpaces2023, vangreevenbroekTradingRegionalOverall2025]. Showcasing many near-optimal solutions can help stakeholders understand the flexibility in planning of systems; if we describe near-optimal spaces geometrically, we can furthermore systematically map out trade-offs.
@@
The space of near-optimal solutions is a convex polytope, being the feasible space of a linear program (LP), but typically consists of too many vertices in too many dimensions to be directly useful. Thus, we follow Grochowicz et al. [grochowiczIntersectingNearoptimalSpaces2023] by approximating a projection onto certain key dimensions of the near-optimal feasible space; the approximation with a limited number of vertices is obtained by repeatedly solving the original model LP with a near-optimality constraint using different, random objective functions.
In this study, the key dimensions we project down to are precisely the variables that participants are allowed to control with sliders: total investment in wind power, solar, heat storage (both overground water tank and geothermal), hydrogen storage (including storage tanks, electrolysers and fuel cells) as well as the total cost of green fuel imports (including ammonia, biogas and pellets). These five dimensions were identified as key technologies for the energy transition in Longyearbyen based on the transition plan, developed by the community council on request by the Norwegian Ministry of Petroleum and Energy (now Ministry of Energy) [grotteEnergiplanLongyearbyenEnergiomstilling2023]. The transition plan evaluated different technological alternatives for electricity and heat supply in Longyearbyen, taking local needs and resource availability into account. Although the transition plan compared the alternatives with the existing diesel generators, we excluded diesel as an MGA dimension and slider (as well as reverting to the old coal power plant) based on the political decisions that Longyearbyen should support and be in line with Norway’s climate targets for 2030 and 2050 [thenorwegianministryofjusticeandpublicsecurityMeldSt262024, grotteEnergiplanLongyearbyenEnergiomstilling2023]. However, the existing diesel generators are still present in the model and do generate electricity. The high operational expenses means that it primarily acts as a back-up, in the absence of (cheaper) alternatives.
@@
We chose a maximum slack level of 125% to allow vastly different system configurations; country-level studies often use lower values instead, with one hindcasting study concluding in historic deviations from the cost-optimum of 9–24% in the UK [trutnevyteDoesCostOptimization2016]). The method presented in Grochowicz et al. [grochowiczIntersectingNearoptimalSpaces2023] leads to an approximation of the near-optimal space by a specified number of boundary vertices — the above paper using 150 optimisations. For the interpolation of metrics to work well, this study is based instead on a total of 56,050 optimisations at a number of different slack levels between 0 and 125%. This way, we obtain near-optimal solutions (for which metrics can be computed) throughout the interior and not only on the boundary of the near-optimal space.
@@
    where $c$ is 1 - the share of VRE, $d$ is 1 - share of imported energy, $e$ is 1 - the share of different technologies present in the system at an installed capacity of at least 1 MW, and $f=\min\{1,heat_{pnorm}\}$ with $heat_{pnorm}$ the normalised capacity of heat storage (a value of $1$, corresponding to $4$ GWh, can cover heat demand for two weeks in the winter, and one month in the summer). In effect, a higher share of VRE, imported energy, as well as a higher complexity of the system increase vulnerability (the value of $V$), whereas higher heat storage reduces vulnerability. The components of the vulnerability metric is based on the transition plan for Longyearbyen [grotteEnergiplanLongyearbyenEnergiomstilling2023], which mention that complexity of the system, dependency on weather, reliance on import and heat storage are all important for the security of supply of the system. The weighting is based on our own conception of the hierarchy among those factors.
@@
Table 2: Assumed land use requirements for technologies included in the modelling (either explicitly in the sliders or implicitly in the model). The land use requirement of different technologies is based on Chen et al. [chenBalancingGHGMitigation2022] and Sasse et al. [sasseRegionalImpactsElectricity2020].
@@
The interface presented to study participants consists of two elements (figure 3): sliders allowing the user to select and tweak an energy system design, and a display of a number of metrics associated with the selected system design (see above). The sliders represent the five key technologies, as explained above. Although participants were only able to change the investment into these five dimensions, the corresponding system designs include both operational decisions as well as investment decisions for other (supporting) technologies (e.g. battery storage) listed in table 1. These five key technologies are the primary influencing factors of the metrics in figure 3b, with the exception of CO2 emissions, which is only based on the electricity generated from the diesel generators (which is not a slider). The sliders as such allow a user to select an energy system design that matches their preferred balance of metrics, even if the selected design contains additional system components.
Both for the sliders in the interface and in the MGA methodology described earlier, we use monetary investment (million NOK) into different technologies as the indicator, as opposed to e.g. electricity generation. This is partly because the five dimensions are different by nature, and do not all align well with electricity generation. For example, the heat infrastructure dimension is related to installing additional heat storage capacity (MWh). Additionally, we want to maintain a focus on the trade-off between costs and other objectives. Although participants clearly showed that they care for more than costs, it remains an important aspect in energy system designs.
The interactive interface is openly available in a GitHub repository (https://github.com/koen-vg/near-opt-interface/) and can be cloned and run on a personal computer. Alternatively, we provide an exemplary use-case for a hypothetical participant in the Supplementary Information.
Importantly, the sliders only allow participants to select feasible system designs at most 125% more expensive than cost-optimal. Formally, the slider values may only encode the coordinates of points within the near-optimal space of the underlaying energy system model. This is achieved by dynamically limiting the selectable ranges of each slider based on the currently chosen design. Specifically, let $x_{1},x_{2},x_{3},x_{4},x_{5}$ represent currently selected slider values, corresponding to the key dimensions described above (wind, solar, green fuel imports, heat infrastructure and hydrogen storage), and let $\mathcal{F}_{\varepsilon}$ by the projection of the near-optimal space down to these five dimensions, with $\varepsilon=125\%$. Then $(x_{1},x_{2},x_{3},x_{4},x_{5})\in\mathcal{F}_{\varepsilon}$. The interface user can change a single variable at a time to select a new system design; without loss of generality, suppose this is $x_{1}$. Allowable new values for $x_{1}$ are exactly those $x_{1}^{*}$ such that $(x_{1}^{*},x_{2},x_{3},x_{4},x_{5})\in\mathcal{F}_{\varepsilon}$. Since $\mathcal{F}_{\varepsilon}$ is convex, this condition is satisfied on a line segment $x_{1}^{*}\in[a,b]$, with
@@
When the user happens to exactly select a system design for which pre-computed metrics were recorded, the interface simply displays those pre-computed values. Since slider values can be changed continuously (up to the limits of pixel and floating point precision) however, it is much more likely that the user selects a system design for which no pre-computed metrics are available. In this case, a heuristic is used to compute approximate values for each of the metrics. The heuristic consists of a linear interpolation among nearby pre-computed metric values. Linear interpolation of metrics has previously been suggested by Lau et al. [lauModellingGenerateContinuous2024], who treat points in a near-optimal space as linear combinations of all computed near-optimal model solutions. While their method works well with a relatively small sample of near-optimal solutions, we apply a near-neighbour filter first in order to further reduce the computational complexity of the heuristic.
@@
Table 3: Demographic characteristics of the sample, the population of Longyearbyen and Norway as a whole00footnotetext: Data for Norway is from 2023 and information of their income was optional, leading to a smaller sample.
@@
Table 5: Exemplary qualitative feedback from the participants. They are referred to in text through their feedback ID.
Feedback ID Quote
R02 “Interesting interlinked scenario. Funny to play around with different parameters and see how complex the energy mix is and consequences are.”
R03 “Well done and gets you thinking.”
R04 “Cannot choose desired energy solution (i.e. 100% return to coal)”
R05 “With respect to the current situation, a coal-fired power plant, which was already in operation, would be the most sensible economic and environmentally friendly solution. Longyearbyen is not equipped for the green transition”
R07 “Look at nuclear power. Safe, space-effective and durable.”
[…]
@@
Author contributions All authors conceived the idea and designed the study. O.V., A.G. and K.vG. did the field work and collected the data. A.G. led the implementation of the near-optimal modelling. K.vG. led the development of the interactive interface. O.V. led the writing of the paper, with considerable input from the other authors.
@@
The field data that support part of the findings of this study are not available owing to the confidentiality of participants’ identities. Longyearbyen is a small settlement, and providing individual data points may risk the identification of particular individuals, even if the data is anonymised. Aggregated data can be made available upon request. The processing of personal data has been assessed and deemed lawful and in compliance with data protection legislation by Sikt – Norwegian Agency for Shared Services in Education and Research (reference number: 188418). The study participants were informed of the processing of personal data, the measures taken to ensure confidentiality and consented to the processing of personal data before participating.
[…]
Declaration of interests The authors declare no competing interests.